The JORDAN & AMANAY Company is in the process of preparing their financial statements for the YEAR ended December 31,2016. The Company has provided you with the following selected account balances from its UNADJUSTED trial balance: DEBIT $40,000 28,000 380,000 CREDIT Cash Accounts Receivable Machinery Accumulated Depreciation Insurance expense Supplies expense Service Revenue Accounts payable Miscellaneous expenses $15,000 24,000 6,000 21,000 40,000 5,000 ADDITIONAL DATA: (a) At the end of the year it was determined that S2000 of supplies were still on hand. When the (b) The employees have earned money for work done during December, but for which they will not be supplies were purchased they were recorded in supplies expense. paid until January of 2017. The amount was $7,000. (c) The Company paid their fire insurance of $24,000 on November 1, 2016 for 6 months in advance. The company recorded the entire amount as insurance expense The machinery is expected to have a useful life of 10 years with an estimated scrap value of $20,000 (d) (e) Customers paid the Company $21,000 in advance on December 1 for services to be performed equally over 3 months-December 2016, January (2017) and February (2017). The company recorded the entire amount as service revenue. REQUIRED: ON THE NEXT PAGE PREPARE THE PROPER ADJUSTING ENTRIES REQUIRED AT 12-31-16. SHOW ALL COMPUTATIONS CLEARLY. USE THE FORMAT PROVIDED. The company adjusts its books YEARL Y!
ADJUSTING ENTRY COMPUTATION DR: CR: DR: CR: DR: CR: DR: CR: DR: CR:
Answer #1
Adjusting Entry Amount Computation
a DR: Supplies 2000 Original/Wrong Entry
           CR: Supplies Expense 2000 DR: Supplies Expense
           CR: Cash
Correct Entry
DR: Supplies
           CR: Cash
b DR: Employee Expense/Salary 7000 No entry passed yet for employees work done in December
           CR: Outstanding Expense 7000
c DR: Prepaid Insurance Expense 16000 Calcualtion for 4 months prepaid rent
           CR: Insurance Expense 16000 24000*4/6 = 16000
d DR: Depreciation Expense 36000 Calculation of Depreciation per year
           CR: Accumulated Depreciation 36000 (380000-20000)/10 = 36000
e DR: Service Revenue 14000 Calculation of Prereceived Income
           CR: Pre received Income 14000 21000*2/3 = 14000